OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that models the values in the table? x -2 0 4 f(x) 4 -6 70 a. y = 3x² + 4x – 6 b. y = 4x² + 3x – 6 c. y = 6x² + 3x – 4 d. y = –4x² – 3x + 6

OpenStudy (anonymous):

if we put x= 0 in all the given equations we get a) y= -6 b) y= -6 c) y= -4 d) y = +6 thus we can directly guess that our required equation is in the upper two equations

OpenStudy (anonymous):

so then it's either going to be A or B

OpenStudy (anonymous):

yes ofcourse now we have to plug in x = 2 and x = 4 in A and B if these function that will give us the desired value will be our answer

OpenStudy (anonymous):

sorry for my bad english

OpenStudy (anonymous):

now lets try x = 2 in A and B

OpenStudy (anonymous):

I'm not getting to work for B

OpenStudy (anonymous):

for x = -2 a) y = 3(4) +4(-2)-6 = 12 - 8-6 = 4-6 = -2 now put x = -2 in b b) y = 4(4)+3(-2) - 6 = 16-6-6 = 10 - 6 = 4 thus our function is B

OpenStudy (anonymous):

Ohhh I see what I did wrong. oops. That makes sense now

OpenStudy (anonymous):

Thanks for your help.

OpenStudy (anonymous):

you are welcome